In 2018, Terrell, Inc., purchases machinery costing $2,528,000. Its 2018 taxable income before considering the Section 179 deduction is $990,000. Assume that Terrell elects not to claim bonus depreciation.

a. Terrell's maximum Section 179 deduction in 2018 is $.

b. The depreciable basis of the equipment is $.
Answer #1

Answer a:

Section 179 maximum limit = $2,500,000

Reduction in section 179 deduction = $25280000 - 25000000 = $28,000

Terrell's maximum Section 179 deduction in 2018 is = $1,000,000 -28000 = $972,000

Terrell's maximum Section 179 deduction in 2018 is = $972,000

Answer b:

The depreciable basis of the equipment is = 2528000 - 972000 = $15,560,000

The depreciable basis of the equipment is = $15,560,000